# The galaxy-IGM connection in thesan: observability and information content of the galaxy-Lyman- cross-correlation at z6

### v1

@@ Appendix D GaLCC evolution fit
We report here the evolution of the amplitude and position of the peak in the GaLCC computed from the thesan-1 simulation. In order to identify them, we smooth the computed GaLCC (using a running mean with window size 2 cMpc/) to remove fluctuations and then identify the maximum of the curve as the peak of the cross-correlation. We have verified that the peak location and height are robust to the size of the window used in the smoothing procedure, and report their evolution in Fig.Ā 23 using diamond and plus symbols, respectively). We separately fit the peak position and height evolution with a function of the form (in analogy to the mean free path evolution found by Worseck etĀ al., 2014):
@@
and report the best-fit evolution with a dashed and dot-dashed line in the Figure. Finally, we report the best-fitting parameter in TableĀ 3.

### v2

@@ Appendix E GaLCC evolution fit
We report here the evolution of the amplitude and position of the peak in the GaLCC computed from the thesan-1 simulation. In order to identify them, we smooth the computed GaLCC (using a running mean with window size 2 cMpc/) to remove fluctuations and then identify the maximum of the curve as the peak of the cross-correlation. We have verified that the peak location and height are robust to the size of the window used in the smoothing procedure, and report their evolution in Fig. 25 using diamond and plus symbols, respectively). We separately fit the peak position and height evolution with a function of the form (in analogy to the mean free path evolution found by Worseck et al., 2014):
@@
and report the best-fit evolution with a dashed and dot-dashed line in the Figure. Finally, we report the best-fitting parameter in Table 3.